Consider the following values for PV of All Benefits, PV of All Costs, and Yearly NPV for year 1-5 of an economic feasibility analysis worksheet, as indicated below:

Yea 1 Year 2 Year 3 Year 4 Year 5
PV of All Benefits 1226 2432 3619 4789 5942
PV of All Costs 3274 4012 4718 5383 6011
Yearly NPV 2047 467 482 504 526
Commulitive NPV ? ? ? ? ?

1- Indicate the value for the Cumulative NPV for year 2?

2- Indicate the value for the Cumulative NPV for year 3?

3-- Indicate the value for the Cumulative NPV for year 4?

4- Indicate the value for the Cumulative NPV for year 5?

5- Indicate the return on investment for this project (percentage)?

6- Describe when will the break-down occur and why?
Answer #1
Year 1 Year 2 Year 3 Year 4 Year 5
PV of All Benefits 1226 2432 3619 4789 5942
PV of All Costs 3274 4012 4718 5383 6011
Yearly NPV 2047 467 482 504 526
Commulative NPV 2047 2514 2996 3500 4026

1- Cumulative NPV for year 2 = 2047+467 = 2514

2- Cumulative NPV for year 3 = 2514+482 = 2996

3- Cumulative NPV for year 4 = 2996+504 = 3500

4- Cumulative NPV for year 5 = 3500+526 = 4026

5- Return on investment = Sum of Yearly NPVs / Sum of PV of Costs = 4026 / (3274+4012+4718+5383+6011) = 17.2 %